OpenStudy (anonymous):

1 more homework question please help!

OpenStudy (anonymous):

\[\log_{4}(\log_{2}16) \]

OpenStudy (knowledge):

consider what is \[\log _{2}16\]

OpenStudy (anonymous):

16=2^x?

OpenStudy (knowledge):

yeah, what is the x

OpenStudy (anonymous):

log(4)

jimthompson5910 (jim_thompson5910):

keep in mind the base is 4

OpenStudy (anonymous):

because of 16?

OpenStudy (knowledge):

x is 4 because 2^4=16

jimthompson5910 (jim_thompson5910):

the outer most log

OpenStudy (knowledge):

now that you know it is four the thing you have to do is evaluate this \[\log _{4}4\]

OpenStudy (knowledge):

what is 4=4^y

OpenStudy (anonymous):

one sec let me catch up

OpenStudy (anonymous):

so 2^log(4) becomes 2^4

OpenStudy (anonymous):

i don't know what 4=4^y is

OpenStudy (knowledge):

what power of four is equal to four

OpenStudy (anonymous):

1

OpenStudy (knowledge):

\[\log _{4}(\log _{2}16)\] \[\log _{4}4\] \[1\]

OpenStudy (knowledge):

so those are the steps to solve this problem

OpenStudy (anonymous):

THANK YOU